OpenStudy (anonymous):

Solve the inequality |x/(x-1)|<2by interpreting the absolute value function as a distance function.

OpenStudy (amistre64):

how do we find the distance between 2 points?

OpenStudy (amistre64):

given y1 = a, and y2 = b d^2 = (a-b)^2

OpenStudy (amistre64):

but they seem to just want you to modify the abs function and not the whole thing |x| = sqrt(x^2)